The credit terms offered to customers for early payment need to be sufficiently lucrative for them to want to pay early, but not so lucrative that the seller is effectively paying an inordinately high interest rate for the use of the money that it is receiving early.

The term structure used for credit terms is to first state the number of days you are giving customers from the invoice date in which to take advantage of the early payment credit terms.

For example, if a customer is supposed to pay within 10 days without any discount, the terms are "net 10 days," whereas if the customer must pay within 10 days to qualify for a 2% discount, the terms are "2/10". To expand upon the last example, if the customer must pay within 10 days to obtain a 2% discount, or can make a normal payment in 30 days, then the terms are stated as "2/10 net 30".

In brief,

                  2% for 20 days

                  y% for 360 days

y = 360 x 2 / 20 = 36 (%)

Meaning that discount of 2% for paying 20 days early equates to an annual interest rate of 36 %.

It is clear that buyers with sufficient cash balances or a readily available line of credit should take advantage of the early payment discounts. However, some buyers are operating with very little cash and have to borrow additional money. These buyers may be wise to forgo the early payment discounts in order to avoid the risk of overdrawing their checking account.

One overdraft fee could be greater than the early payment discount, which makes the discount pointless to the buyer.

Problem setting Fill the gaps.

Credit
Terms


Explanation

Discount
interest

Overdraft Interest rate

Enjoy a discountas a buyer?

Net 30

Pay in 30 days

None

2%

Yes

1/10 Net 30

Take 1% discount if pay in 10 days, otherwise pay in 30 days

18%

20%

No

2/10 Net 30

Take 2% discount if pay in 10 days, otherwise pay in 30 days

36%

20%

Yes

1/10 Net 60

Take 1% discount if pay in 10 days, otherwise pay in 60 days

10%

2/10 Net 60

Take 2% discount if pay in 10 days, otherwise pay in 60 days

14,4%

10%

3/10 Net 30

Take 3% discount if pay in 10 days, otherwise pay in 30 days

10%

3/10 Net 60

Take 3% discount if pay in 10 days, otherwise pay in 60 days

18%
Answer #1

Please see the table below:

I have also added a column "How it has been calculated" so that you understand the mathematics behind the same.

Credit Terms Explanation Discount Interest Overdraft Interest rate Enjoy a discount as a buyer? How it has been calculated?
Net 30 Pay in 30 days None 2% Yes (This answer should be not applicable or N.A. as there is no discount under this scheme)
1/10 Net 30 Take 1% discount if pay in 10 days, otherwise pay in 30 days 18% 20% No
2/10 Net 30 Take 2% discount if pay in 10 days, otherwise pay in 30 days 36% 20% Yes
1/10 Net 60 Take 1% discount if pay in 10 days, otherwise pay in 60 days 7.20% 10% No For 60 - 10 = 50 days, discount offered is 1%; hence discount for 360 days should be = 1% / 50 x 360 = 7.20%
2/10 Net 60 Take 2% discount if pay in 10 days, otherwise pay in 60 days 14.40% 10% Yes For 60 - 10 = 50 days, discount offered is 2%; hence discount for 360 days should be = 2% / 50 x 360 = 14.40%
3/10 Net 30 Take 3% discount if pay in 10 days, otherwise pay in 30 days 54.00% 10% Yes For 30 - 10 = 20 days, discount offered is 3%; hence discount for 360 days should be = 3% / 20 x 360 = 54.00%
3/10 Net 60 Take 3% discount if pay in 10 days, otherwise pay in 60 days 21.60% 18% Yes For 60 - 10 = 50 days, discount offered is 3%; hence discount for 360 days should be = 3% / 50 x 360 = 21.60%